# Spectral Mode Enhancement in Coherent-harmonic Dual-comb Spectroscopy Enables Exceeding 300-fold averaging-time reduction

### v3

@@ II.2 H-H-DCS
As illustrated in Fig. 3, C-H-DCS utilizes two H-OFCs with pulse repetition rates of $4f_{\mathrm{ms}}$ and $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$. In the time domain, similar to a C-DCS with the same repetition rate difference of $4\Delta f_{\mathrm{s}}$, this architecture generates 4 centerbursts within a time period of $1/\Delta f_{\mathrm{s}}$. Due to this repetition rate difference, the phase-inverted pulses in the two pulse trains gradually slip relative to each other, resulting in distinct in-pulse-pair phase configurations: $[\pi,0,0,0]-[\pi,0,0,0]$, $[0,\pi,0,0]-[\pi,0,0,0]$, $[0,0,\pi,0]-[\pi,0,0,0]$, and $[0,0,0,\pi]-[\pi,0,0,0]$, respectively, in the time domain regions corresponding to the four centerbursts. Consequently, the $\phi_{TPM}$ in Eq. 1, is constantly 0 for the first centerburst; while for the other three centerbursts, the values are periodic $[\pi,\pi,0,0],[\pi,0,\pi,0]$, and $[\pi,0,0,\pi]$, respectively. These time-varying phases (related to the pulse pair index, i.e., time) increase the oscillation frequency of the interference electric field, shifting the 2nd to 4th centerbursts to higher-order components, which can be filtered out by an electronic low-pass filter in practical implementation. In fact, these high-order components correspond to spectral copies generated by the heterodyning of teeth from comb 1 with non-nearest teeth from comb 2 in the spectrum, as detailed in Supplement 1.
@@
In the frequency domain, two H-OFCs with 2-fold mode amplitude enhancement and periodic multilevel phase spectra of $[-\pi,-\pi,-\pi,0]$, beat together to yield an RF comb. Since the amplitude of a heterodyne beat note is the product of the amplitudes of the two beating components, and the phase is the phase difference between them, the modes of this RF comb therefore exhibit a 4-fold enhancement in amplitude and possess a uniform phase, which precisely corresponds to the single fundamental centerburst with a 4-fold increase in amplitude in the time domain.
Thus, this architecture leads to a 4-fold SNR gain. For a more general case, this number 4 corresponds to m, and is equivalent to an SNR gain brought by coherent averaging spanning a total time of $m^{2}$-fold.
Interestingly, the aforementioned IGM characteristics can be changed (the SNR gain is not changed) when the two TPM signals use the quadratic form described in Eq. 3 with the same m but different s values. The generated RF comb can exhibit quadratic periodic multilevel phase profiles of $-\pi\frac{p_{1}-p_{2}}{m}j^{2}$. According to the temporal self-imaging effect 38, 30, the fundamental centerburst on the IGM will then be redistributed, similar to that in C-H-DCS.
@@ III Experiment
Fig. 4: Sketch of experimental setup. Rb clock: rubidium clock; OFC: optical frequency comb; AWG: arbitrary waveform generator; RFA: RF amplifiers; EOPM/EOIM: electro-optical phase/intensity modulator; OC: optical coupler; PD: photodetector; LPF: low-pass filter.
Fig. 5: Comparison of interferograms (IGMs) and signal-to-noise ratio (SNR) for the three architectures. (a) IGMs of the C-DCS and (b) its spectrum (partial). The centerburst repetition rate is 1 kHz, matching the spectral line spacing, as expected for conventional DCS. The spectrum exhibits a 20 kHz amplitude periodicity, with prominent lines attributed to the non-ideal extinction ratio of the EOIMs 39. (c) IGMs of the C-H-DCS and (d) its spectrum. The centerburst repetition rate is observed at 20 kHz, a 20th harmonic of its line spacing of 1 kHz, corresponding to the average mode amplitude in the spectrum increasing to $>$-60 dB. (e) IGMs of the H-H-DCS and (f) its spectrum. The centerburst repetition rate and the line spacing are both observed to be 1 kHz. The amplitude of the centerbursts is significantly increased to a value approaching 20 mVpp, corresponding to the spectral mode amplitude increased to close to -50 dB. (g) Evolution of time-domain SNR for the H-H-DCS and the C-DCS with the number of averages. On the log-log scale, the linear fitting models of the two datasets are given by fit 1 = 0.4997(4)x + 2.497(1) and fit 2 = 0.5075(4)x + 1.238(1). The slopes, approximating 0.5, indicate that the SNR is proportional to the square root of the number of averages, consistent with the principle of coherent averaging. The time-domain SNR is defined as the ratio of the centerburst amplitude to the standard deviation of a segment away from the centerburst 40.
@@
Moreover, the two OFCs operated at repetition rates of 250 MHz and 249.98 MHz with a $\Delta f_{\mathrm{rep}}$ = 20 kHz, and carrier-envelope offset frequencies ($f_{\mathrm{ceo}}$) of 20 MHz and 33.5 MHz, respectively. To ensure long-term stability of the system, these $f_{\mathrm{ceo}}$ and $f_{\mathrm{rep}}$ were simply phase-locked to a rubidium atomic clock, with which the AWG was also timing synchronized (variable delay lines or RF phase shifters can be employed to align the optical pulses and modulation signals). Additionally, the high degree of mutual coherence essential for DCS was retrieved through the optical referencing and post-treatment techniques 19, 42, 43, which can be regarded as an original digital version of the adaptive scheme 44.
For the DCS experiments, an $\mathrm{H^{13}C^{14}N}$ gas cell (Wavelength References, HCN-13-10), with a nominal pressure of approximately 10 Torr and a path length of 16.5 cm, was selected as a representative sample. The two modulated combs were mixed and subsequently split into measurement and reference beams using a 50/50 coupler. The measurement beam traversed the $\mathrm{H^{13}C^{14}N}$ gas cell before detection by a biased detector (Thorlabs, DET01CFC), while the reference beam was directly detected. This reference channel is crucial for eliminating comb tooth fluctuations arising from the uneven gain caused by the error in $\phi_{TPM}$. IGMs from both detectors were low-pass filtered (6 MHz @3 dB) and simultaneously digitized by an oscilloscope at a sampling rate of 31.25 MS/s. The optical power incident on the detector in the measurement beam was $11.9~{}\mu W$ , $119~{}\mu W$, and $223~{}\mu W$ for the C-DCS, the C-H-DCS, and the H-H-DCS, respectively, limited by filter bandwidth and amplifier limitations (Supplementary Fig. S1).
@@ IV Results and discussion
Fig. 6: Example $\mathrm{H^{13}C^{14}N}$ spectra relative to the center frequency of 193.5449 THz (1548.9555 nm) at a spectral resolution of 12.5 MHz, along with the corresponding weighted Voigt profile fits. (a) Recorded in 1.999 s (1999 averages) in the C-DCS. (b) Recorded in 1.999 s in the C-H-DCS. (c) Recorded in 5 ms (5 averages) in the H-H-DCS. (d) Recorded in 1.999 s in the H-H-DCS. The standard deviations of the four fitting residuals were calculated, with the residual noise in (d) is reduced by a factor of $\sim$20 compared to that in (a).
@@ IV.1 Features of Interferograms
Figure 5(a–f) displays the IGMs generated in the three architectures and their corresponding spectra after post treatment. As anticipated, the centerburst repetition rate in the C-H-DCS is 20 times higher than that of the C-DCS, while maintaining the same spectral mode spacing and achieving an average enhancement of the mode amplitude by approximately 7 dB ($\mathrm{10lg\sqrt{20}}$).
@@
As for the H-H-DCS, the centerburst amplitude is approximately 20 times greater than that in the C-DCS. This corresponds to an average enhancement of the mode amplitude by approximately 13 dB ($\mathrm{10lg20}$) in the spectrum. The noise floor in the spectrum also increases by about 2 dB, potentially due to higher relative intensity noise. Furthermore, in addition to the primary centerbursts, the IGMs of the H-H-DCS also show weak centerbursts with a 1/20 ms interval, which is attributed to imperfect transfer of the oscillation frequency due to errors in the driving signal levels. The mode superposition of these weak centerbursts caused the unflatness on the RF comb, i.e., an amplitude modulation with a period of 20 kHz, as shown in Figs. 5(e) and (f). Details regarding the higher-order components are presented in Supplementary Fig. S3.
A potential misconception might arise that the amplified centerburst amplitude necessitates a larger detector dynamic range. In fact, as mentioned in the principle section, this is a superposition resulting from the convolution of the additional pulse pairs due to low-pass filtering.
IV.2 Time-domain Signal-to-noise Radio
The evolution of the time-domain SNR for the H-H-DCS and the C-DCS with respect to the number of averages (or measurement time) is calculated in Fig. 5(g). The x-axis intercepts of the two linear fits exhibit a difference of 2.5575, indicating that this new architecture reduces the required number of averages or measurement time to achieve the same SNR by a factor of 361 (1/361 of the original). The y-axis intercepts show a difference of 1.259, indicating an approximate 18-fold enhancement in SNR for the same number of averages. These findings are consistent with the expected performance for a setup with m = 20.
IV.3 Transmission Spectrum
@@
The enhancement in time-domain SNR is manifested in the transmission spectrum. Fig. 6 presents the acquired transmission spectra around the P7 absorption line of $\mathrm{H^{13}C^{14}N}$ using the three architectures, with a spectral point spacing of 12.5 MHz. Both architectures, particularly the H-H-DCS, demonstrate significantly reduced residual levels within the same measurement time compared to the C-DCS. Furthermore, the standard deviations of the fitting residual observed in a 5-ms measurement using H-H-DCS are comparable to those in a $\sim$ 2-s measurement using C-DCS.
@@ IV.4 Discussion
The proof-of-concept results presented in this work can be further improved by combining with ultrahigh-repetition-rate comb sources, or when probing with finer spectral resolution. For example, when combined with 22-GHz microresonator combs 45 and operating at a spectral resolution of 500 MHz, compared to conventional DCS with the same resolution, an exceeding 1500-fold averaging-time reduction is expected. Furthermore, integrated lithium niobate electro-optic modulators operating at CMOS-compatible voltages may facilitate this implementation 46.
In wavelength regions where mature EOPMs are not readily available, this approach may can be achieved using nonlinear means. For example, under second-harmonic generation 47, a temporal pulse-phase profile of $[0,0,0,\pi/2]$ is anticipated to transform to the desired $[0,0,0,\pi]$, and a profile of $\pi\frac{s}{2m}k^{2}$ is anticipated to transform to the desired $\pi\frac{s}{m}k^{2}$.
For field-deployed DCS, the mitigation in coherence time inherently offers significant advantages. Further possible combination with ultra-low-noise optical frequency combs, such as the compact monolithic solid-state design 48 that provides a $\sim$ 100 Hz linewidth in a quasi-free-running state, complex and sophisticated locking systems may be greatly simplified. Moreover, the decoupling of the centerburst repetition rate and the $\Delta f_{\mathrm{rep}}$ (i.e. $\Delta f_{\mathrm{s}}$ ) also overcomes the limitation on the noise tracking bandwidth of post-treatment algorithms in free-running DCS, which was previously limited to half the IGM repetition rate (i.e. $\Delta f_{\mathrm{rep}}/2$) 43, 49, a restriction that can be roughly attributed to the fact that in conventional DCS, only a single centerburst within each IGM provides effective sampling.
Finally, it is worth emphasizing that the essence of this spectral mode enhancement lies in enabling the optical frequency comb or combs to output multiple special harmonic pulse trains with the same carrier-envelope offset frequency, but different carrier-envelope phases (non-linear slip), thereby enhancing the majority of the mode sets rather than enhancing a few and suppressing others. The scheme demonstrated in this paper, which employs TPM based on EOPMs outside the cavity, may not be the only method to achieve this SNR enhancement. For example, in some parametrically driven Kerr soliton microresonators, two soliton states with opposite phases owing to the $Z_{2}$-symmetry have been observed 50, and a soliton sequence with random phase bits of $[0,0,\pi,0]$ has been demonstrated 51.

### v4

@@ II.3 PH-PH-DCS
As illustrated in Fig. 3, PH-PH-DCS utilizes two PH-OFCs with pulse repetition rates of $4f_{\mathrm{ms}}$ and $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$. In the time domain, similar to a C-DCS with the same repetition rate difference of $4\Delta f_{\mathrm{s}}$, this architecture generates 4 centerbursts within a time period of $1/\Delta f_{\mathrm{s}}$. Due to this repetition rate difference, the phase-inverted pulses in the two pulse trains gradually slip relative to each other, resulting in distinct in-pulse-pair phase configurations: $\pi,0,0,0]-[\pi,0,0,0]$, $\pi,0,0]-[\pi,0,0,0], [0,0,$$0,\pi,0]-[\pi,0,0,0$], and $0,0,0,\pi]-[\pi,0,0,0]$, respectively, in the time domain regions corresponding to the four centerbursts. Consequently, the $\phi_{TPM}$ in Eq. 3, is constantly 0 for the first centerburst; while for the other three centerbursts, the values are periodic $\pi,\pi,0,0],[\pi,0,\pi,0]$, and $\pi,0,0,\pi]$, respectively. These time-varying phases (related to the pulse pair index, i.e., time) increase the oscillation frequency of the interference electric field, shifting the 2nd to 4th centerbursts to higher-order components, which can be filtered out by an electronic low-pass filter in practical implementation. In fact, these high-order components correspond to spectral copies generated by the heterodyning of teeth from comb 1 with non-nearest teeth from comb 2 in the spectrum, as detailed in Supplement 1.
@@
In the frequency domain, two PH-OFCs with 2-fold mode amplitude enhancement compared to the C-OFC and periodic multilevel phase spectra of $\pi,\pi,\pi,0]$, beat together to yield an RF comb. Since the amplitude of a heterodyne beat note is the product of the amplitudes of the two beating components, and the phase is the phase difference between them, the modes of this RF comb therefore exhibit a 4-fold enhancement in amplitude and possess a uniform phase, which precisely corresponds to the single fundamental centerburst with a 4-fold enhancement in amplitude in the time domain. For a more general case, the amplitude enhancement is m-fold.
A potential misconception might arise that the amplified centerburst amplitude necessitates a larger detector dynamic range. In fact, this is a superposition resulting from the convolution of the additional pulse pairs due to low-pass filtering. Interestingly, the aforementioned IGM characteristics can be changed when the two TPM signals use the quadratic form described in Eq. 2 with the same m but different s values. The generated RF comb can exhibit quadratic periodic multilevel phase profiles of $\pi\frac{p_{1}-p_{2}}{m}j^{2}$. According to the temporal self-imaging effect 37, 27, the fundamental centerburst on the IGM will then be redistributed to m/2 copies, similar to that in C-PH-DCS. A similar effect is expected when using two PRBS signals with different patterns. This will help to alleviate the dynamic range pressure on components within the detection chain other than the detector.
@@ III Experiment
The two OFCs operated at repetition rates of 250 MHz and 249.98 MHz with a $\Delta f_{\mathrm{rep}}$ = 20 kHz, and carrier-envelope offset frequencies ($f_{\mathrm{ceo}}$) of 20 MHz and 33.5 MHz, respectively. To ensure long-term stability of the system, these $f_{\mathrm{ceo}}$ and $f_{\mathrm{rep}}$ were simply phase-locked to a rubidium atomic clock, with which the AWG was also timing synchronized. Variable delay lines or RF phase shifters can be employed to align the optical pulses and modulation signals. The high degree of mutual coherence essential for DCS was retrieved through the optical referencing and post-treatment techniques40. A continuous wave (CW) laser is used as an intermediate oscillator 41 and the self-correction algorithm based on cross-correlation function 42, 43 are implemented.
Furthermore, operating at spectral bin spacing of 12.5 MHz leads to a limited available spectral bandwidth. According to the DCS non-aliasing condition 14, the available bandwidth satisfies $\Delta\nu\leq f_{\mathrm{ms}}^{2}/{2\Delta f_{\mathrm{s}}}$, which yields 78.125 GHz (0.626 nm at 1550 nm) here for $f_{\mathrm{ms}}$ = 12.5 MHz and $\Delta f_{\mathrm{s}}$ = 1 kHz. Therefore, optical band-pass filters (BPFs) were employed to narrow the spectrum, specifically two 1.3-nm @ 25 dB filters (BPF 1) and a 0.3-nm @ 12 dB filter (BPF 2). Because most of the optical power was lost during filtering, two EDFAs were used to partially compensate. The optical power incident on the detector in the measurement beam was $1.9~\mu W, 11$$19~\mu W$, and $223~\mu W$ for the C-DCS, the C-PH-DCS, and the PH-PH-DCS.
For the DCS measurement, an $\mathrm{H^{13}C^{14}N}$ gas cell (Wavelength References, HCN-13-10) with a nominal pressure of approximately 10 Torr and a path length of 16.5 cm was selected as a representative sample. In addition to the measurement beam, a reference beam was directly detected for baseline normalization. Both beams were detected by two InGaAs biased photodetectors (Thorlabs DET01CFC, 1 GHz bandwidth, no internal amplification). The photocurrent was low-pass filtered (6 MHz @ 3 dB) and then converted to a voltage signal by terminating the detector directly into the 50 $\Omega$ input of a Tektronix MSO54 oscilloscope. The signal was then digitized at a sampling rate of 31.25 MS/s with 12-bit resolution.
@@ IV Results and discussion
Fig. 5: Comparison of interferograms (IGMs) and spectrum for the three architectures. (a) IGMs of the C-DCS and (b) its spectrum (partial). The centerburst repetition rate is 1 kHz, matching the spectral line spacing, as expected for conventional DCS. The spectrum exhibits a 20 kHz amplitude periodicity, with prominent lines attributed to the non-ideal extinction ratio of the EOIMs 38. (c) IGMs of the C-PH-DCS and (d) its spectrum. The centerburst repetition rate is observed at 20 kHz, a 20th harmonic of its line spacing of 1 kHz, corresponding to the average mode amplitude in the spectrum increasing to >>-60 dB. (e) IGMs of the PH-PH-DCS and (f) its spectrum. The centerburst repetition rate and the line spacing are both observed to be 1 kHz. The amplitude of the centerbursts is significantly increased to a value approaching 20 mVpp, corresponding to the spectral mode amplitude increased to close to -50 dB.
@@ IV.1 Features of Interferograms
Figure 5(a–f) displays the raw IGMs generated in the three architectures and their corresponding spectra after post treatment. As anticipated, the centerburst repetition rate in the C-PH-DCS is 20 times higher than that of the C-DCS, while maintaining the same spectral mode spacing and achieving an average enhancement of the mode amplitude by approximately 7 dB ($\mathrm{10lg\sqrt{20}}$).
@@
As for the PH-PH-DCS, the centerburst amplitude is approximately 20 times greater than that in the C-DCS. This corresponds to an average enhancement of the mode amplitude by approximately 13 dB ($\mathrm{10lg20}$) in the spectrum. The noise floor in the spectrum also increases by about 2 dB. Furthermore, in addition to the primary centerbursts, the IGMs of the PH-PH-DCS also show weak centerbursts with a 1/20 ms interval. This is attributed to an error in the modulation signal level, which prevented the original centerbursts at these adjacent locations from being completely transferred to high-frequency components (Fig. 3(a)), leaving a small residue at the fundamental oscillation frequency. The mode superposition of these weak centerbursts caused the unflatness on the RF comb, i.e., an amplitude modulation with a period of 20 kHz, as shown in Figs. 5(e) and (f). Details regarding the higher-order components are discussed in Supplement 1.
IV.2 Signal-to-Noise Ratio and Noise Floor
The power spectral densities (PSDs) of the noise sections of the three DCS are shown in Fig. 6. In a single 1-ms measurement, none of them reached the shot-noise limit. This shortfall is primarily due to the optical power loss introduced by bandpass filtering. The optical power increased by a factor of about $\sim$17x from C-DCS to PH-PH-DCS, whereas the noise power only increased by a factor of 1.74x (2.4 dB). This discrepancy is likely due to quantization noise or relative intensity noise (RIN).
The evolution of the time-domain SNR for the PH-PH-DCS and the C-DCS as a function of the number of averages (or measurement time) is shown in Fig. 6 (c). The x-axis separation of 2.463 between fit 2 and fit 1 at the same SNR level indicates that this new architecture reduces the averaging time requirement by a factor of 290, i.e., down to 1/290 of the original time. The y-axis intercepts differ by 1.222, corresponding to a 16.7-fold SNR enhancement at the same number of averages. These observations agree well with the expected performance for a multiplexing factor of m = 20 in approximately constant-noise-limited region.
Nevertheless, if the system operates in the shot-noise limited region 18, the SNR enhancement and averaging time reduction are expected to decrease to factors of 4 and 17, respectively. More generally, for an arbitrary multiplexing factor m, the SNR scales as $\sqrt{m}$ and the averaging time reduces by a factor of m.
Fig. 6: (a) Noise sections of interferogram (IGM) used to calculate PSDs. The main and small centerbursts and their surrounding regions containing the free induction decay are excluded. (b) PSDs of the noise sections of the three DCS. Three colored dashed lines represent the theoretical shot noise limits for the corresponding DCS in a single measurement (1 ms), calculated from the optical power and detector response. The shaded region from 3 to 6 MHz corresponds to the location of the interferogram spectrum. Within this region, in a single measurement, the noise power of C-PH-DCS (green) and PH-PH-DCS (red) increased by 0.5 dB and 2.5 dB, respectively, compared to C-DCS (blue). In a $\sim$2 s measurement (1999 averages), the increase was 0.75 dB and 2.4 dB, respectively. (c) Evolution of time-domain SNR for the PH-PH-DCS and the C-DCS with the number of averages. On the log-log scale, the linear fitting models of the two datasets are given by Fit 1 = 0.4851(4)x + 2.475(2) and Fit 2 = 0.4962(4)x + 1.253(1). The slopes, approximating 0.5, indicate that the SNR is proportional to the square root of the number of averages, consistent with the principle of coherent averaging. The time-domain SNR is defined as the ratio of the centerburst amplitude to the standard deviation of some segments away from the centerburst 44.
Fig. 7: Example $\mathrm{H^{13}C^{14}N}$ spectra relative to the center frequency of 193.5449 THz (1548.9555 nm) at a spectral resolution of 12.5 MHz, along with the corresponding weighted Voigt profile fits. (a) Recorded in 1.999 s (1999 averages) in the C-DCS. (b) Recorded in 1.999 s in the C-PH-DCS. (c) Recorded in 1.999 s in the PH-PH-DCS.
IV.3 Baseline Normalization and Transmission Spectrum
As mentioned, the baselines of these spectra were normalized by the second reference channel. This is because the actual modulated phase levels deviate from the ideal values, leading to mode-to-mode variations in the spectrum, as shown in Figs. 5 (d) and (f). Such variations are also commonly observed in DCS setups based on electro-optic (EO) combs 13 and certain microcombs. These high-frequency baseline fluctuations are often difficult to correct using conventional self-fitting methods 45.
Nevertheless, in PPHE-DCS, such fluctuations exhibit a periodicity of $m$-times mode spacing. Utilizing this periodic characteristic, a spectrum of PPHE-DCS can be deinterleaved into m normal spectra, which can then be fitted and processed individually as in C-DCS, without relying on the second channel. A detailed demonstration is provided in Supplement 1.
@@ IV.3 Baseline Normalization and Transmission Spectrum
The enhancement in time-domain SNR is evident in the transmission spectrum. Fig. 7 presents the transmission spectra acquired around the P7 absorption line of $\mathrm{H^{13}C^{14}N}$ with the three architectures, with a spectral point spacing of 12.5 MHz. Both architectures, particularly the PH-PH-DCS, demonstrate significantly lower residual levels within the same measurement time compared to the C-DCS, with the residual noise reduced by a factor of $\sim$20.
It is worth noting that the three weighted Voigt profile fits show slight differences in line depth, we attribute this to the differences in SNR and tracking bandwith of the cross-correlation function 46.
@@ IV.4 Discussion
We expect this concept can be integrated with ultra-high-repetition-rate OFC source, such as microcombs 47. In that case, this PPHE-DCS can achieve a higher SNR enhancement or averaging time reduction, and become applicable to typical use cases requiring spectral sampling spacings from hundreds of MHz to the GHz range 21.
Finally, it is worth discussing that the PH-OFC share an essential similarity with harmonic mode-locking: the pulse train exhibits correlation between every mth pulse 32, whereas neighboring pulses of which are not fully correlated. Under this condition, the spectrum contains an interleaving of m supermodes, resulting in an effective mode spacing of $f_{\mathrm{rep}}/m$ rather than $f_{\mathrm{rep}}$. These additional supermodes, apart from the dominant one, are conceptually similar to the ”densified” or ”self-imaging” modes here. Therefore, We speculate that the extra-cavity-TPM-modulating scheme, may not be the only viable approach to realize this concept. For instance, conventional harmonically mode-locked lasers aim for strong supermode suppression 48. In contrast, if low supermode suppression could be achieved such that the m supermodes have comparable intensities, it might provide a suitable source for PPHE-DCS. Moreover, in some parametrically driven Kerr soliton microresonators, two soliton states with opposite phases owing to the $Z_{2}$-symmetry have been observed 49, and a soliton sequence with random phase bits of $0,0,\pi,0]$ has been demonstrated 50.